OpenStudy (anonymous):

What is the inverse of y = 3x + 9?

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

To find the inverse of a function. Change places of x and y. Then, solve for y again.

OpenStudy (anonymous):

y=-3 so plug it in now ?

OpenStudy (anonymous):

x=-4

OpenStudy (ipwnbunnies):

What? Did you find the inverse function?

OpenStudy (anonymous):

i solved it

OpenStudy (ipwnbunnies):

By switching places, I mean like this: x = 3y + 9 Then solve for y. That'll be the inverse of the function.

OpenStudy (ipwnbunnies):

Oh. If the original function has a point at (-4,-3) then the inverse has a point at (-3,-4)

OpenStudy (anonymous):

thats what i did, and i got y=-3, then i switched them back and pluged y in to solve for x

OpenStudy (anonymous):

thats not one of the answers tho, the're fractions but none of them have 4 in them

OpenStudy (ipwnbunnies):

I'm confused now. The instructions just say to find the inverse.

OpenStudy (ipwnbunnies):

We're not inputting anything in. We're just finding the inverse function.

OpenStudy (anonymous):

A. y=1/3x-3 B. y = 9x + 3 C. y=1/3x+9 D. y=-1/3x-3

OpenStudy (anonymous):

i think its d

OpenStudy (ipwnbunnies):

Just find it like the way I showed you. Original function: y = 3x + 9 Switch variables: x = 3y + 9 Can you rearrange this to solve for y?

OpenStudy (anonymous):

so its d

OpenStudy (ipwnbunnies):

No.

OpenStudy (anonymous):

thats the only one with a negative

OpenStudy (ipwnbunnies):

x - 9 = 3y; divide both sides by 3 x/3 - 3 = y

OpenStudy (anonymous):

so its a

OpenStudy (ipwnbunnies):

Yes

OpenStudy (anonymous):

thank you, your awesome !